2- if 60 g of 1% hydrocortisone is mixed with 80 g of 2.5% hydrocortisone, what is the percent W/W of hydrocortisone in the final mixture?

10 - How would 500 ml bag of 3% 50/50 be prepared using 3%sodium chloride, sodium acetate 2 meq/ml, and sterile water for injection.

9- How many g of the drug are required to dispense one pint of 1 in 25 solutions

7- How many g of g of sodium chloride should be used to prepare 1.5 l of normal saline? How many ml of a 23.4%of sodium chloride will be used to make this solution? What base solution will be used?

6- Bupivacaine 0.125% in 400 ml 0.9% sodium chloride premix bag are out of stock. How would you prepare a bag using bupivacaine 0.25%

5- How many ml of sodium chloride 4meq/ml are needed to make 1l D10/0.45% sodium chloride using 1L bag of dextrose 10%
Answer #1

solution:

2.

Step 1:

Amount of cortisone in 60gm 1%

Amount of cortisone = 60 /100 X 1 = 0.6g

Step 2:

Amount of cortisone in 80gm is 2.5%

Amount of cortisone = 80 / 100 X 2.5= 2g

Step 3:

Final mixture = 100 X (0.6+ 2) / 60 + 80

= 100 X 2.6 / 140

= 1.85% W/W.

amount of cortisone in the final mixture is 1.85% W/W